# Finite-time blowup in a five- and higher-dimensional chemotaxis model with indirect signal production

### v2

@@ 4. Energy bounded below by sublinear dissipation rate
for brevity. Using $w=-y+v-f$ and $w_{r}=-y_{r}+v_{r}-f_{r}$, we evaluate for $\eta\in(0,R)$
$\int_{0}^{\eta}wv_{r}r^{\alpha}\;\mathrm{d}r$ $=\int_{0}^{\eta}(-y+v-f)v_{r}r^{\alpha}\;\mathrm{d}r$ (4.14)
$=-\int_{0}^{\eta}(v_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r+\frac{1}{2}\int_{0}^{\eta}(v^{2})_{r}r^{\alpha}\;\mathrm{d}r-\int_{0}^{\eta}fv_{r}r^{\alpha}\;\mathrm{d}r$
$=-\frac{1}{2}v_{r}^{2}(\eta)\eta^{\alpha}+\frac{1}{2}v^{2}(\eta)\eta^{\alpha}-(n-1)\int_{0}^{\eta}v^{2}r^{\alpha-1}\;\mathrm{d}r-\int_{0}^{\eta}fv_{r}r^{\alpha}\;\mathrm{d}r$
@@
$-\int_{0}^{\eta}w_{r}yr^{\alpha}\;\mathrm{d}r=-\int_{0}^{\eta}(-y_{r}+v_{r}-f_{r})yr^{\alpha}\;\mathrm{d}r$ (4.15)
$=\int_{0}^{\eta}yy_{r}r^{\alpha}\;\mathrm{d}r-\int_{0}^{\eta}(v_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r+\int_{0}^{\eta}f_{r}yr^{\alpha}\;\mathrm{d}r$
$=\frac{1}{2}y^{2}(\eta)\eta^{\alpha}-(n-1)\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r-\frac{1}{2}v_{r}^{2}(\eta)\eta^{\alpha}+\int_{0}^{\eta}f_{r}yr^{\alpha}\;\mathrm{d}r.$
Collecting (4.12)–(4.15) and multiplying by $\eta^{1-n}$, we get
$\quad\frac{1}{2}y^{2}(\eta)\eta^{n-1}+w_{r}(\eta)v_{r}(\eta)\eta^{n-1}-\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r$ (4.16)
$=\frac{1}{2}v^{2}(\eta)\eta^{n-1}-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}fv_{r}r^{\alpha}\;\mathrm{d}r$
$\quad-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}f_{r}yr^{\alpha}\;\mathrm{d}r-\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}v^{2}r^{\alpha-1}\;\mathrm{d}r$
@@
$\leq\frac{1}{2}v^{2}(\eta)\eta^{n-1}-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r+\omega_{n}^{-1}\|f\|_{L^{2}(B_{\eta})}\|\nabla v\|_{L^{2}(B_{\eta})}$
$\quad+\omega_{n}^{-1}\|\nabla f\|_{L^{2}(B_{\eta})}\|\Delta v\|_{L^{2}(B_{\eta})}$
@@
for all $\eta\in(0,R)$, where the last inequality follows from dropping the last term and Hölder inequality. Integrating (4.16) over $(0,\rho)\subset(0,R)$, we have
$\frac{1}{2}\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}w_{r}v_{r}\eta^{n-1}\;\mathrm{d}\eta-\int_{0}^{\rho}\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r\;\mathrm{d}\eta$ (4.17)
$\leq\frac{1}{2}\int_{0}^{\rho}v^{2}\eta^{n-1}\;\mathrm{d}\eta-\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r\;\mathrm{d}\eta$
@@
$\quad+\omega_{n}^{-1}\rho\|\nabla f\|_{L^{2}(B_{\rho})}\|\Delta v\|_{L^{2}(B_{\rho})}+\omega_{n}^{-1}\rho\|f\|_{L^{2}(B_{\rho})}\|\nabla v\|_{L^{2}(B_{\rho})}.$
@@
We calculate the second term on the left of (4.17) by integration by parts, the identity $w=-y+v-f$ from (4.1) and Hölder inequality
$\int_{0}^{\rho}w_{r}v_{r}\eta^{n-1}\;\mathrm{d}\eta=wv_{r}\rho^{n-1}-\int_{0}^{\rho}wy\eta^{n-1}\;\mathrm{d}\eta$ (4.18)
$=wv_{r}\rho^{n-1}+\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta-\int_{0}^{\rho}vy\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}fy\eta^{n-1}\;\mathrm{d}\eta$
$=wv_{r}\rho^{n-1}-vv_{r}\rho^{n-1}+\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}v_{r}^{2}\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}fy\eta^{n-1}\;\mathrm{d}\eta$
$\geq\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta+\omega_{n}^{-1}\|\nabla v\|_{L^{2}(B_{\rho})}^{2}-\omega_{n}^{-1}\|\Delta v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}+(w-v)v_{r}\rho^{n-1}$
for all $\rho\in(0,R)$. By Fubini theorem, we compute the last term on the left of (4.17)
$-\int_{0}^{\rho}\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r\;\mathrm{d}\eta=-(n-1)\int_{0}^{\rho}\int_{r}^{\rho}\frac{\;\mathrm{d}\eta}{\eta^{n-1}}y^{2}r^{\alpha-1}\;\mathrm{d}r$ (4.19)
$=-\frac{n-1}{n-2}\int_{0}^{\rho}y^{2}r^{n-1}\left(1-\frac{r^{n-2}}{\rho^{n-2}}\right)\;\mathrm{d}r\geq-\frac{n-1}{n-2}\int_{0}^{\rho}y^{2}r^{n-1}\;\mathrm{d}r.$
@@
Thank to $uv_{r}=u_{r}-g\sqrt{u}$ as given by (4.2), the second term on the right of (4.17) can be controlled accordingly
$-\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r\;\mathrm{d}\eta=-\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}(u_{r}-g\sqrt{u})r^{\alpha}\;\mathrm{d}r\;\mathrm{d}\eta$ (4.20)
$\leq-\int_{0}^{\rho}u\eta^{n-1}\;\mathrm{d}\eta+\alpha\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}ur^{\alpha-1}\;\mathrm{d}r\;\mathrm{d}\eta+\int_{0}^{\rho}\int_{0}^{\eta}|g\sqrt{u}|r^{n-1}\;\mathrm{d}r\;\mathrm{d}\eta$
@@
$\leq-\int_{0}^{\rho}u\eta^{n-1}\;\mathrm{d}\eta+\frac{\alpha}{n-2}\int_{0}^{\rho}u\eta^{n-1}\left(1-\frac{\eta^{n-2}}{\rho^{n-2}}\right)\;\mathrm{d}\eta+\frac{\sqrt{m}\rho}{\omega_{n}}\|g\|_{L^{2}(B_{\rho})}$
$\leq\frac{nm}{\omega_{n}(n-2)}+\frac{\sqrt{m}\rho}{\omega_{n}}\|g\|_{L^{2}(B_{\rho})}.$
Substituting (4.18)–(4.20) into (4.17) we get
@@
$\left(\frac{3}{2}-\frac{n-1}{n-2}\right)\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+\|\nabla v\|_{L^{2}(B_{\rho})}^{2}$ (4.21)
$\leq\rho\left(\|\nabla f\|_{L^{2}(B_{\rho})}\|\Delta v\|_{L^{2}(B_{\rho})}+\|\nabla v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}\right)+\|\Delta v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}$
$\quad 2m+\sqrt{m}\rho\|g\|_{L^{2}(B_{\rho})}+\|v\|_{L^{2}(B_{\rho})}^{2}+\omega_{n}(v-w)v_{r}\rho^{n-1}.$
The pointwise restrictions (4.5) imply
@@
thank to $n\geq 5$, we plug five estimates above into (4.21), which becomes (4.11) by putting $C_{0}:=\max\{\omega_{n},12+R^{2}\}$. ∎
We shall eliminate the zeroth-order terms $\|f\|_{L^{2}(\Omega)}^{2}$ and $\|v\|_{L^{2}(\Omega)}^{2}$ left in Lemma 4.2 and Lemma 4.3 without treatment.
@@
Let $n\geq 5$. Then there exist $C>0$ and $C(\kappa)>0$ such that
$-\frac{1}{24}\mathcal{F}$ $\leq 12\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+CB^{4/(n+2)}\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\sqrt{m}\rho\|g\|_{L^{2}(\Omega)}$
$\quad+C(\kappa)(B^{2}+m^{2}+M^{2}+1)\rho^{n-2\kappa}$
holds for all $(u,v)\in\mathcal{S}$ and $\rho\in(0,R)$, where $\mathcal{F}$ and $\mathcal{D}$ are specified as in (4.3).
@@
Lemma 4.1 and 4.2 entail there exists $C(\kappa)>0$ such that
$-\mathcal{F}$ $\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(\Omega)}^{2}+C(\kappa)(Bm+BM+1)\rho^{2-\kappa}$
holds for all $\rho\in(0,R)$. Furthermore, Lemma 4.3 warrants
$-\frac{1}{24}\mathcal{F}$ $\leq 12\rho^{2}\|\nabla f\|_{L^{2}(B_{\rho})}^{2}+C_{0}\|f\|_{L^{2}(B_{\rho})}^{2}+\sqrt{m}\rho\|g\|_{L^{2}(B_{\rho})}$ (4.22)
$\quad+\frac{9}{8}\|v\|_{L^{2}(B_{\rho})}^{2}-\frac{3}{4}\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+C(\kappa)(B^{2}+Bm+BM+m+1)\rho^{n-2\kappa}$
@@
holds for all $\psi\in W^{1,2}(\Omega)$. Recalling (4.4) we thus have
$\|f\|_{L^{2}(\Omega)}^{2}\leq C_{gn}B^{4/(n+2)}\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+C_{gn}B^{2},$ (4.23)
@@
$\|v\|_{L^{2}(B_{\rho})}^{2}$ $\leq C_{gn}M^{4/(n+2)}\|\nabla v\|_{L^{2}(\Omega)}^{2n/(n+2)}+C_{gn}M^{2}$ (4.24)
$\leq\frac{1}{2}\|\nabla v\|_{L^{2}(\Omega)}^{2}+2^{n}C_{gn}^{(n+2)/2}M^{2}+C_{gn}M^{2}$
for all $\rho\in(0,R)$ by Young inequality. Invoking (4.5), we estimate
$\|\nabla v\|^{2}_{L^{2}(\Omega)}$ $=\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+\|\nabla v\|^{2}_{L^{2}(\Omega\setminus B_{\rho})}$ (4.25)
$\leq\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+B^{2}\||x|^{1-\kappa}\|_{L^{2}(\Omega\setminus B_{\rho})}^{2}$
$\leq\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+\omega_{n}B^{2}R^{2}\rho^{n-2\kappa}$
for all $\rho\in(0,R)$. Inserting (4.23)–(4.25) into (4.22) yields
$-\frac{1}{24}\mathcal{F}$ $\leq 12\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+C_{0}C_{gn}B^{4/(n+2)}\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\sqrt{m}\rho\|g\|_{L^{2}(\Omega)}$
$\quad+C(\kappa)(B^{2}+Bm+M^{2}+m+1)\rho^{n-2\kappa}$
for all $\rho\in(0,R)$. We finish the proof. ∎
Here we shall reproduce the elementary proof which was given by Winkler [47] to bound the energy $\mathcal{F}$ from below by a sublinear power of the dissipation rate $\mathcal{D}$.
@@
Let $n\geq 5$. Then there exist $\theta\in(0,1)$ and $C(m,M,B,\kappa)>0$ such that
$-\mathcal{F}\leq C(m,M,B,\kappa)(\mathcal{D}^{\theta}+1)$ (4.26)
holds for all $(u,v)\in\mathcal{S}$, where $\mathcal{F}$ and $\mathcal{D}$ are specified as in (4.3).
@@
$\rho:=\min\left\{\frac{R}{2},\|\nabla f\|_{L^{2}(\Omega)}^{-\frac{2n}{(n+2)(2\kappa-n)}}\right\},$
@@
$\theta:=\begin{cases}\frac{n}{n+2}>1-\frac{2n}{(n+2)(2\kappa-n)},&\text{if }\kappa\in(n-2,n),\\ 1-\frac{2n}{(n+2)(2\kappa-n)}\geq\frac{n}{n+2},&\text{if }\kappa\geq n.\end{cases}$

### v3

@@ 4 Energy bounded below by sublinear dissipation rate
for brevity. Using w=y+vfw=-y+v-f and $w_{r}=-y_{r}+v_{r}-f_{r}$, we evaluate for $\eta\in(0,R)$
$\begin{split}&\int_{0}^{\eta}wv_{r}r^{\alpha}\;\mathrm{d}r=\int_{0}^{\eta}(-y+v-f)v_{r}r^{\alpha}\;\mathrm{d}r\\ &=-\int_{0}^{\eta}(v_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r+\frac{1}{2}\int_{0}^{\eta}(v^{2})_{r}r^{\alpha}\;\mathrm{d}r-\int_{0}^{\eta}fv_{r}r^{\alpha}\;\mathrm{d}r\\ &=-\frac{1}{2}v_{r}^{2}(\eta)\eta^{\alpha}+\frac{1}{2}v^{2}(\eta)\eta^{\alpha}-(n-1)\int_{0}^{\eta}v^{2}r^{\alpha-1}\;\mathrm{d}r-\int_{0}^{\eta}fv_{r}r^{\alpha}\;\mathrm{d}r\end{split}$ (35)
@@
$\begin{split}&-\int_{0}^{\eta}w_{r}yr^{\alpha}\;\mathrm{d}r=-\int_{0}^{\eta}(-y_{r}+v_{r}-f_{r})yr^{\alpha}\;\mathrm{d}r\\ &=\int_{0}^{\eta}yy_{r}r^{\alpha}\;\mathrm{d}r-\int_{0}^{\eta}(v_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r+\int_{0}^{\eta}f_{r}yr^{\alpha}\;\mathrm{d}r\\ &=\frac{1}{2}y^{2}(\eta)\eta^{\alpha}-(n-1)\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r-\frac{1}{2}v_{r}^{2}(\eta)\eta^{\alpha}+\int_{0}^{\eta}f_{r}yr^{\alpha}\;\mathrm{d}r.\end{split}$ (36)
Collecting (33)–(36) and multiplying by $\eta^{1-n}$, we get
@@
$\begin{split}&\frac{1}{2}y^{2}(\eta)\eta^{n-1}+w_{r}(\eta)v_{r}(\eta)\eta^{n-1}-\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r\\ =&\frac{1}{2}v^{2}(\eta)\eta^{n-1}-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}fv_{r}r^{\alpha}\;\mathrm{d}r\\ &\quad-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}f_{r}yr^{\alpha}\;\mathrm{d}r-\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}v^{2}r^{\alpha-1}\;\mathrm{d}r\\ \leq&\frac{1}{2}v^{2}(\eta)\eta^{n-1}-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r+\omega_{n}^{-1}\|f\|_{L^{2}(B_{\eta})}\|\nabla v\|_{L^{2}(B_{\eta})}\\ &\quad+\omega_{n}^{-1}\|\nabla f\|_{L^{2}(B_{\eta})}\|\Delta v\|_{L^{2}(B_{\eta})}\end{split}$ (37)
@@
for all $\eta\in(0,R)$, where the last inequality follows from dropping the last term and Hölder inequality. Integrating (37) over $(0,\rho)\subset(0,R)$, we have
@@
$\begin{split}&\frac{1}{2}\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}w_{r}v_{r}\eta^{n-1}\;\mathrm{d}\eta-\int_{0}^{\rho}\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r\;\mathrm{d}\eta\\ \leq&\frac{1}{2}\int_{0}^{\rho}v^{2}\eta^{n-1}\;\mathrm{d}\eta-\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r\;\mathrm{d}\eta\\ &\quad+\omega_{n}^{-1}\rho\|\nabla f\|_{L^{2}(B_{\rho})}\|\Delta v\|_{L^{2}(B_{\rho})}+\omega_{n}^{-1}\rho\|f\|_{L^{2}(B_{\rho})}\|\nabla v\|_{L^{2}(B_{\rho})}.\end{split}$ (38)
@@
We calculate the second term on the left of (38) by integration by parts, the identity w=y+vfw=-y+v-f from (21) and Hölder inequality
$\begin{split}&\int_{0}^{\rho}w_{r}v_{r}\eta^{n-1}\;\mathrm{d}\eta=wv_{r}\rho^{n-1}-\int_{0}^{\rho}wy\eta^{n-1}\;\mathrm{d}\eta\\ &=wv_{r}\rho^{n-1}+\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta-\int_{0}^{\rho}vy\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}fy\eta^{n-1}\;\mathrm{d}\eta\\ &=wv_{r}\rho^{n-1}-vv_{r}\rho^{n-1}+\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}v_{r}^{2}\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}fy\eta^{n-1}\;\mathrm{d}\eta\\ &\geq\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta+\omega_{n}^{-1}\|\nabla v\|_{L^{2}(B_{\rho})}^{2}-\omega_{n}^{-1}\|\Delta v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}+(w-v)v_{r}\rho^{n-1}\end{split}$ (39)
for all $\rho\in(0,R)$. By Fubini theorem, we compute the last term on the left of (38)
$\begin{split}&-\int_{0}^{\rho}\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r\;\mathrm{d}\eta=-(n-1)\int_{0}^{\rho}\int_{r}^{\rho}\frac{\;\mathrm{d}\eta}{\eta^{n-1}}y^{2}r^{\alpha-1}\;\mathrm{d}r\\ =&-\frac{n-1}{n-2}\int_{0}^{\rho}y^{2}r^{n-1}\left(1-\frac{r^{n-2}}{\rho^{n-2}}\right)\;\mathrm{d}r\geq-\frac{n-1}{n-2}\int_{0}^{\rho}y^{2}r^{n-1}\;\mathrm{d}r.\end{split}$ (40)
@@
Thanks to $uv_{r}=u_{r}-g\sqrt{u}$ given by (22), the second term on the right of (38) can be controlled accordingly
@@
$\begin{split}&-\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r\;\mathrm{d}\eta=-\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}(u_{r}-g\sqrt{u})r^{\alpha}\;\mathrm{d}r\;\mathrm{d}\eta\\ \leq&-\int_{0}^{\rho}u\eta^{n-1}\;\mathrm{d}\eta+\alpha\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}ur^{\alpha-1}\;\mathrm{d}r\;\mathrm{d}\eta+\int_{0}^{\rho}\int_{0}^{\eta}|g\sqrt{u}|r^{n-1}\;\mathrm{d}r\;\mathrm{d}\eta\\ \leq&-\int_{0}^{\rho}u\eta^{n-1}\;\mathrm{d}\eta+\frac{\alpha}{n-2}\int_{0}^{\rho}u\eta^{n-1}\left(1-\frac{\eta^{n-2}}{\rho^{n-2}}\right)\;\mathrm{d}\eta+\frac{\sqrt{m}\rho}{\omega_{n}}\|g\|_{L^{2}(B_{\rho})}\\ \leq&\ \frac{nm}{\omega_{n}(n-2)}+\frac{\sqrt{m}\rho}{\omega_{n}}\|g\|_{L^{2}(B_{\rho})}.\end{split}$ (41)
Substituting (39)–(41) into (38) we get
@@
$\begin{split}&\left(\frac{3}{2}-\frac{n-1}{n-2}\right)\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+\|\nabla v\|_{L^{2}(B_{\rho})}^{2}\\ &\leq\rho\left(\|\nabla f\|_{L^{2}(B_{\rho})}\|\Delta v\|_{L^{2}(B_{\rho})}+\|\nabla v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}\right)+\|\Delta v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}\\ &\quad 2m+\sqrt{m}\rho\|g\|_{L^{2}(B_{\rho})}+\|v\|_{L^{2}(B_{\rho})}^{2}+\omega_{n}(v-w)v_{r}\rho^{n-1}.\end{split}$ (42)
The pointwise restrictions (26) imply
@@
thanks to $n\geq 5$, we plug five estimates above into (42), which becomes (32) by putting $C_{0}:=\max\{\omega_{n},12+R^{2}\}$. ∎
We shall eliminate the quadratic zeroth-order terms $\|f\|_{L^{2}(\Omega)}^{2}$ and $\|v\|_{L^{2}(\Omega)}^{2}$ left in Lemma 4.2 and Lemma 4.3 without treatment.
@@
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ be a ball of radius $R>0$ with some $n\geq 5$. Then there exists a constant $C_{4}>0$ such that
[…]
$\quad+C_{4}(R^{2\kappa-n}+1)(B^{2}+m^{2}+M^{2}+1)\rho^{n-2\kappa}$
holds for all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$ and $\rho\in(0,R)$, where $\mathcal{F}$ is given in (24).
@@
Lemma 4.1 and 4.2 entail there exists $C_{1}>0$ such that
-$\mathcal{F}$ $\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(B_{\rho})}^{2}+C_{1}(Bm+BM)\rho^{2-\kappa}+C_{1}$
holds for all $\rho\in(0,R)$. Furthermore, Lemma 4.3 warrants the existence of $C_{3}>0$ satisfying
$\begin{split}&-\frac{1}{24}\mathcal{F}\leq 12\rho^{2}\|\nabla f\|_{L^{2}(B_{\rho})}^{2}+C_{3}\|f\|_{L^{2}(B_{\rho})}^{2}+\sqrt{m}\rho\|g\|_{L^{2}(B_{\rho})}\\ &+\frac{9}{8}\|v\|_{L^{2}(B_{\rho})}^{2}-\frac{3}{4}\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+C_{3}B^{2}\rho^{n-2\kappa}+C_{3}(Bm+BM)\rho^{2-\kappa}+2m+C_{1}\end{split}$ (43)
@@
holds for all $\psi\in W^{1,2}(\Omega)$. Recalling (25) we thus have
$\|f\|_{L^{2}(\Omega)}^{2}\leq C_{gn}B^{4/(n+2)}\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+C_{gn}B^{2},$ (44)
@@
$\begin{split}\|v\|_{L^{2}(B_{\rho})}^{2}&\leq C_{gn}M^{4/(n+2)}\|\nabla v\|_{L^{2}(\Omega)}^{2n/(n+2)}+C_{gn}M^{2}\\ &\leq\frac{1}{2}\|\nabla v\|_{L^{2}(\Omega)}^{2}+2^{n}C_{gn}^{(n+2)/2}M^{2}+C_{gn}M^{2}\end{split}$ (45)
for all $\rho\in(0,R)$ by Young inequality. Invoking (26), we estimate
$\begin{split}\|\nabla v\|^{2}_{L^{2}(\Omega)}&=\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+\|\nabla v\|^{2}_{L^{2}(\Omega\setminus B_{\rho})}\\ &\leq\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+B^{2}\||x|^{1-\kappa}\|_{L^{2}(\Omega\setminus B_{\rho})}^{2}\\ &\leq\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+\omega_{n}B^{2}R^{2}\rho^{n-2\kappa}\end{split}$ (46)
for all $\rho\in(0,R)$. Inserting (44)–(46) into (43) yields
$\begin{split}-\frac{1}{24}\mathcal{F}&\leq 12\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+C_{3}C_{gn}B^{4/(n+2)}\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\sqrt{m}\rho\|g\|_{L^{2}(\Omega)}+C(\rho)\end{split}$
for all $\rho\in(0,R)$, where
$C(\rho)$ $:=C_{3}C_{gn}B^{2}+2^{n+1}C_{gn}^{(n+2)/2}M^{2}+2C_{gn}M^{2}+\omega_{n}B^{2}R^{2}\rho^{n-2\kappa}$
$\quad+C_{3}B^{2}\rho^{n-2\kappa}+C_{3}(Bm+BM)\rho^{2-\kappa}+2m+C_{1}$
$\>\leq(C_{3}C_{gn}+2^{n+1}C_{gn}^{(n+2)/2}+2C_{gn}+1)(B^{2}+M^{2}+m^{2}+1)$
$\quad+(\omega_{n}R^{2}+C_{3}+C_{3}R^{\kappa-n+2}+C_{1}R^{2\kappa-n})(B^{2}+M^{2}+m^{2}+1)\rho^{n-2\kappa}.$
Taking
$\begin{split}C_{4}&:=C_{3}C_{gn}+2^{n+1}C_{gn}^{(n+2)/2}+2C_{gn}+1+\omega_{n}+2C_{3}+C_{1},\end{split}$
we finish the proof. ∎
Here we shall choose $\rho$ appropriately, following the strategy of Winkler [6], to establish a lower bound of the energy $\mathcal{F}$ by a sublinear power of the dissipation rate $\mathcal{D}$.
@@
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ be a ball of radius $R>0$ with some $n\geq 5$. Then there exist $\theta\in(0,1)$ and $C(m,M,B,\kappa)>0$ such that
$\mathcal{F}\leq C(m,M,B,\kappa)(\mathcal{D}^{\theta}+1)$ (47)
holds for all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$, where $\mathcal{F}$ and $\mathcal{D}$ are specified as in (24).
@@
$\rho:=\min\left\{\frac{R}{2},\|\nabla f\|_{L^{2}(\Omega)}^{-\frac{n}{(n+2)(2\kappa-n)}}\right\},$
@@
$\theta:=\begin{cases}\frac{n}{n+2}>1-\frac{n}{(n+2)(2\kappa-n)}>0,&\text{if }\kappa\in\big(n-2,\frac{3n}{4}\big),\\ 1-\frac{n}{(n+2)(2\kappa-n)}\geq\frac{n}{n+2},&\text{if }\kappa\in\big[\frac{3n}{4},\infty\big).\end{cases}$